OpenStudy (anonymous):

Solve the equation please: 3 log x+1 = 6-log x

OpenStudy (mrcoolguy):

x = 5/4

OpenStudy (anonymous):

3 log x+1=6-log x 3 logx+ log x=6-1 4 log x=5 \[\log x^4=5\] \[x^4=e^5,x=\left( e^5 \right)^{\frac{ 1 }{ 4 }}=e ^{\frac{ 5 }{ 4 }}\]